OpenStudy (dtan5457):

Someone refresh me on this. How do I write a quadratic equation knowing that the sum of roots are 6 and product of roots are -27?

OpenStudy (dtan5457):

@SolomonZelman

OpenStudy (dtan5457):

@satellite73

OpenStudy (anonymous):

ax^2-(sum of roots)x+(product of roots)=0

OpenStudy (anonymous):

think about what you do when you factor

OpenStudy (anonymous):

also what @MARC_ said, but make the leading coefficient 1

OpenStudy (dtan5457):

so x^2-(-27)x+(6)=0?

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

product of the roots is the constant

OpenStudy (anonymous):

you got it backwards

OpenStudy (anonymous):

yup

OpenStudy (dtan5457):

oh, so switch numbers so you get x^2+6x-27?

OpenStudy (dtan5457):

as the final equation?

OpenStudy (anonymous):

\[x^2-6x-24=0\]

OpenStudy (anonymous):

ur sign is wrong @dtan5457

OpenStudy (anonymous):

@satellite73 -27, not -24 :)

OpenStudy (dtan5457):

-(sum of roots)x wouldn't that be -6x?

OpenStudy (anonymous):

it suppose be x^2-6x-27

OpenStudy (dtan5457):

i didn't write -6x did i

OpenStudy (anonymous):

yup

OpenStudy (dtan5457):

so ax^2-(sum of roots)x+(product of roots)=0 is the formula for any question like this?

OpenStudy (anonymous):

yes @dtan5457